Home CSU, Chico Disney A Home Thrift books iCloud ter 1 ICA#30 Cherokee Inc is a merchandiser that provided the following Information Humber of unit sold 28,888 Sellire price per unit Uariable sellire expense per unit Uariable administrative expense per unit Total fixed sellirg expense Total fixed administ rative expense inning merc hand ise inventory Endire merchandise inventory Merchandise purchases $ 48,800 $ 33,000 $ 24,600 44,808 130,00 okRequired: 1. Prepare a tradtional Income statement 2 Prepare a contribution format Income statement. int Complete this question by entering your answers in the tabs below. Prepare a traditional income statement. Cherokee, Inc Traditional Income Statement Sales Cost of goods sold Gross margin 600,000 160,000 440,000 Selling and administrative expenses Selling expenses Pdministrative expenses Selling expenses Administrative expenses $80,000 40,000 40,000 30,000 190,000 $ 250,000 Net operating income BequE
Answer #1

2) Contribution Format Income Statement :- (Amount in $)

Cherokee Inc.
Contribution format Income Statement
Sales (20,000*30) (a) 600,000
Less: Variable Costs
Cost of Goods Sold (b) 160,000
Selling Expense(20,000*4) (c) 80,000
Administrative Expense(20,000*2) (d) 40,000
Total Variable Costs (e=b+c+d) (280,000)
Contribution Margin(f=a-e) 320,000
Less: Fixed Cost
Selling Expense (g) 40,000
Administrative Expense(h) 30,000
Total Fixed Cost (i=g+h) (70,000)
Net Operating Income(f-i) 250,000

Working Notes:

(b) Cost of Goods Sold:- Beginning Merchandise Inventory + Merchandise Purchases - Ending Merchandise Inventory

=24,000+ 180,000-44,000 =160,000